hi5656:

PLEASE HELP Solve for x: one over four (4x + 15) = 24

hi5656:

:/

jhonyy9:

\[\frac{ 1 }{ 4}\left( 4X +15 \right) = 24\] so like a first step multiplie both sides by 4 and what you get in this way ?